A farmer has 1,110 acres of land on which he grows corn, wheat, and soybeans. It costs $45 per acre to grow corn, $60 to grow wheat, and $50 to grow soybeans. Because of market demand the farmer will grow twice as many acres of wheat as of corn. He has allocated $58,950 for the cost of growing his crops. How many acres of each crop should he plant?
Answer by richwmiller(17219)

c+w+s=1100,
45c+60w+50s=58950,
w=2c
c = 790/3, s = 310, w = 1580/3
c=263.33, s=310, w=526.67